Standard deviating of the breaking strength of 10 cables tested by company was 815kg.Find 95%,and 99%confidence limit for the deviation of all cables produced by the company.
1
Expert's answer
2019-11-11T10:51:00-0500

Standard deviating of the breaking strength of 10 cables tested by company was 815kg with standard deviation 45 kg.Find 95%,and 99%confidence limit for the deviation of all cables produced by the company.

We need to construct the 95% confidence interval for the population mean "\mu"

Since the population's "\sigma" is not known the formula uses the T-distribution with "n-1" degrees of freedom:


"CI=(\bar{X}-t_c\times{s \over \sqrt{n}},\bar{X}+t_c\times{s \over \sqrt{n}})"

Given that

"\bar{X}=815, s=45, df=n-1=10-1=9"

For "\alpha=0.05"


"t_c=z_ {1-\alpha/2;n-1}=2.262"

"CI=(\bar{X}-t_c\times{s \over \sqrt{n}},\bar{X}+t_c\times{s \over \sqrt{n}})="

"=(815-2.262{45 \over \sqrt{10}},815+2.262\times{45 \over \sqrt{10}})="

"=(782.809,847.191)"

We need to construct the 99% confidence interval for the population mean "\mu"

For "\alpha=0.01"


"t_c=z_ {1-\alpha/2;n-1}=3.250"

"CI=(\bar{X}-t_c\times{s \over \sqrt{n}},\bar{X}+t_c\times{s \over \sqrt{n}})="

"=(815-3.250{45 \over \sqrt{10}},815+3.250\times{45 \over \sqrt{10}})="

"=(768.754,861.246)"